# FlexRT - A fast and flexible cosmological radiative transfer code for reionization studies I: Code validation

### v1

@@ 1 Introduction
Looking forward, the deepest insights into reionization will come from combining these rich and complimentary data sets. A key problem confronting the field is how to maximize the information contained in them. Together, they are broadly sensitive to not only the distribution of intergalactic hydrogen during reionization, but also the structures and histories of the HI-ionizing radiation background and IGM temperature. Simulating reionization accurately at this level of detail requires solving, or at least approximating, the radiative transfer (RT) equation, which encodes the physics linking the ionizing sources to the state of the IGM. But solving the RT equation in this context is a formidable computational challenge for several reasons. The first is the high dimensionality of the problem. The RT equation has $7$ dimensions ($3$ spatial, $2$ angular, $1$ frequency, and $1$ time), all of which are important for reionization. Moreover, reionization is a massively multi-scale problem. The intrinsic ionizing properties of galaxies are set by astrophysics on parsec scales and below: star formation and feedback processes, AGN activity, as well as the structures of the interstellar and circumgalactic media. The small-scale clumpiness of the IGM sets its recombination rate and self-shielding properties, which determine the number of photons that the sources must produce to complete reionization [32, 33, 34]. The clumping scale may be as low as a kpc. Finally, simulations in representative volumes ($10$s-$100$s of cMpc on a side) may also contain millions of ionizing sources clustering on $\sim 10$ cMpc scales. The distribution and sizes of ionized regions depends on the source clustering as well as the small-scale clumpiness of the IGM [e.g. 35]. In principle, a fully self-consistent reionization simulation would therefore require at least $8-9$ orders of magnitude in dynamic range.
@@
On the other hand, the fully numerical approach is well-suited for studying in detail the state and physics of the reionizing IGM. In recent years, this approach has generally moved in the direction of fully coupled radiative hydrodynamics simulations, many of which also simulate the galaxies, including star formation, AGN, and feedback processes [e.g. 49, 51, 55, 56, 52]. With this comes the great advantage that the simulations can now leverage the constraints provided by direct observations of the galaxy population. But the state-of-the-art simulations can cost millions of core hours per run, allowing for only a handful of simulations to be run per year, which precludes parameter space studies. The limited number of such simulations available also makes it difficult to disentangle the effects of approximations and differing sub-grid modeling choices, which are required in even the most state-of-the-art simulations [e.g. 57, 58]. It is also worth noting that the moment-based RT methods that are most widely used today start from an approximate form of the RT equation. Hence, these methods also do not converge to the exact RT solution in any limit of the parameters related to numerical implementation (e.g. number of resolution elements).
@@ 2.1.1 Moment-Based RT methods
The high dimensionality of the RT equation makes it one of the most computationally intensive equations to solve numerically, even in relatively simple cases. A common approach to reducing the dimensionality is to take moments of the RT equation in the angular dimensions, and then truncate the hierarchy at the dipole level with an ansatz for the quadrupole moment, i.e. the Eddington tensor. The two most prominent ansatzes in use for reionization codes today are the M1 closure [61] and the optically thin Eddington tensor [OTVET; 62]. The equations governing the first two moments – the radiation energy density and flux – take on the form of conservation laws, analogous to the Euler equations of fluid dynamics, for which very efficient numerical solvers have been developed. Such codes are also relatively straightforward to parallelize across many CPUs because of their fluid-like description of the radiation field. [^2] Indeed, the M1 closure has the particularly convenient property that the Eddington tensor becomes a local function of radiation energy density and flux. Owing to its computational efficiency, the moment approach has become the most widely adopted type of RT algorithm used in almost all of the recent radiation hydrodynamics simulations of reionization, many of which include galaxy formation physics [49, 50, 55, 51, 52]. Indeed, without the moment-based RT implementations, the unprecedented scales and resolutions achieved by these simulations would not be possible.
The accuracy of moment-based codes has been quantified for a suite of idealized test problems in the Cosmological Radiative Transfer Comparison Project [59, 63], and in other works [e.g. 48, 64]. To our knowledge, however, there has been no quantitative study on the inaccuracies introduced by the fluid-like approximation to the radiation field in the cosmological reionization problem. There is good motivation for such a study. Ref. [65] considered several idealized RT problems that are relevant for reionization and found that moment-based methods tend to: (1) over-ionize dense, self-shielding absorption systems, which can result in a factor of $\sim 2$ under-prediction of the photoionization rate of intergalactic gas for a given emissivity; and (2) err in the structure of the ionizing radiation background on scales smaller than the mean free path (MFP). For instance, they found that the M1 closure significantly underestimates the fluctuations on these scales. This effect is visualized nicely in Figure 5 of Ref. [23], who noted that RT simulations run with Aton broadly exhibited a more uniform ionizing background compared to models generated with EX-CITE, which is essentially a ray tracing algorithm.
@@ 3.1 Ray creation and propagation
In traditional time-independent ray tracing schemes, rays would be traced away from the source until they either terminate in a neutral region or all their photons are absorbed by the IGM [76, 46, 23]. In FlexRT, “rays” are treated instead like photon packets that travel a distance $\Delta x_{\rm RT}$, the RT cell size, during each time step. This approach, inspired by Monte-Carlo codes, naturally lends itself to the finite speed of light and “localizes” the rays, making FlexRT easier to parallelize (see §2.1.3). As we will see, this approach makes our procedure for merging of rays particularly straightforward. To avoid numerical artifacts introduced by the HealPix structure, during each time step we randomly rotate the HealPix spheres used to cast rays around each source by a randomly chosen set of direction angles, $\alpha$, $\beta$, and $\gamma$. Each ray retains knowledge of these angles, so that the original HealPix unit vectors associated with it can be recovered when needed. We find that this procedure eliminates discernible structural artifacts even for simulations with just a single ionizing source.
@@ 3.3 Ray merging
Figure 1 shows a 2D representation of the steps in our merging scheme. Panel A shows an RT cell containing a set of rays. The lengths of the rays represent how many photons they contain, and rays with the same colors/line styles point in similar directions. In panel B, we throw down a randomly rotated “pizza” (analogous to a HealPix sphere in 3D) and assign each ray to the slice it occupies. In panel C, rays within the same slice are merged. Note that two of the slices are empty and one of them has only a single ray, so no merging occurs in those slices. In the slice with the red rays, the longest ray had enough photons to be exempt from merging, and the other two were combined. Panel D shows the $x-y$ positions of the merged rays, which are the photon-weighted averaged positions of the constituent rays (see again panel A).
Figure 1: 2D representation of how ray merging works in FlexRT. Panel A shows a cell containing several rays. The lengths of the arrows represent the number of photons in each ray, and rays with same color/line style point in similar directions. In panel B, we throw down a randomly rotated “pizza” (analogous to a HealPix sphere in 3D) and group rays by the “slice” they occupy. We merge rays occupying the same slice in panel C, excluding the longest (red) ray, since it has enough photons to be exempt. Merged rays point along the bisector of the slice they occupy (i.e. the HealPix pixel direction). In panel D, we show that the positions of merged rays are the photon-weighted average of their constituents.
@@ 4.1 Direct solution of the chemistry equations (“Standard Mode”)
In this mode, we assume that all small-scale fluctuations in the HI number density ($n_{\rm HI}$) are fully resolved by our simulation. That is, $n_{\rm HI}$ does not fluctuate below the scale of the RT cell, $\Delta x_{\rm cell}$. Under this assumption, we can write an initial guess for the HI photo-ionization rate $\Gamma_{\rm HI}$ at the beginning of an RT time step of length $\Delta t=\Delta x_{\rm cell}/c$ as:
@@
where $T$ is temperature, $C_{\rm HI}(T)$ is the collisional ionization coefficient of HI, $\alpha(T)$ is the recombination coefficient of HII, $n_{e}$ is the electron number density and $n_{\rm HII}$ is the HII number density. Updating Eq. 4.2 using Eq. 4.1 would ignore the fact that $n_{\rm HI}$ and $\Gamma_{\rm HI}$ may evolve during $\Delta t$. Following Ref. [46], we account for this by solving for the time-dependence of the ionized fraction $x_{i}$ during $\Delta t$ assuming that the recombination rate rate remains constant (Eq. 12-14 of Ref. [46]),
@@
We iterate Eq. 4.3-4.7 several times to achieve convergence. Ref. [46] pointed out that this procedure gives the exact solution for $n_{\rm HI}$, no matter how large $\Delta t$ is, if recombinations are neglected. Ignoring recombinations is a reasonable approximation inside ionization fronts (I-fronts), where $n_{\rm HI}$ evolves most quickly and this procedure is most important.
@@
In the first term on the RHS, $\mathcal{H}$ and $\Lambda$, are the net radiative heating and cooling rates, respectively, $n_{\rm tot}$ is the total number density of gas particles (neutrals, ions, and electrons), and $k_{\rm B}$ is Boltzmann’s constant. We include heating by photo-ionizations as well as cooling by recombinations, collisional ionizations and excitations, free-free interactions, and Compton scattering off the CMB. The second term, wherein $H(z)$ is the Hubble parameter, accounts for adiabatic cooling by cosmic expansion. The third term accounts for changes in temperature from an increase or decrease in the number of particles in the gas. The last term accounts for heating/cooling by compression/expansion, and is proportional to the change in the density contrast $\Delta$.
@@ 4.2 General treatment (“Generalized Opacity Mode”)
FlexRT’s second mode is built around relaxing the assumption that small-scale fluctuations in $n_{\rm HI}$ are fully resolved. Relaxing this assumption allows for each cell to have some un-resolved spatial distribution of $n_{\rm HI}$ inside it that would not be accounted for in Standard Mode. In this case, the equations in the previous section become ill-defined, since they rely on assuming a single value for $n_{\rm HI}$ in each cell. A commonly-employed method to correct for the effect of un-resolved density fluctuations is to boost the recombination rate (and thus, the residual $n_{\rm HI}$ in photo-ionization equilibrium) in each cell by a sub-grid clumping factor [e.g. 38]. However, this is not the most general approach, since it assumes that the ionizing opacity is determined uniquely by the recombination rate (e.g. Eq. 17 of Ref. [78]). This is not true in the presence of dense, self-shielding structures that harbor neutral gas, wherein the assumption of photo-ionization equilibrium may not hold.
@@
$\tau_{\nu}\equiv\kappa_{\nu}\Delta s\equiv\Delta x_{\rm cell}/\lambda_{\nu}$ (4.9)
@@
Our final task is to provide a general expression for $\Gamma_{\rm HI}$ given knowledge of $\kappa_{\nu}$. Indeed, if $\Delta x_{\rm cell}$ were assumed to always be $<<\lambda_{\nu}$, Eq. 4.10 would be sufficient. However, as already mentioned, a key feature of FlexRT is that it should admit the use of large (up to several Mpc) RT cells without substantial loss of accuracy. As such, we must allow for scenarios in which $\lambda_{\nu}\lesssim\Delta x_{\rm cell}$. We can do this as follows. First, we assume (without loss of generality, as we will see) that all rays $j$ during the time step $\Delta t$ are incident on one face of cell $i$ traveling in the direction normal to that face. We can write the contribution of ray $j$ to $\Gamma_{\rm HI}$ at a distance $x_{j}$ into the cell as
@@
where $A_{\rm cell}\equiv\Delta x_{\rm cell}^{2}$ and we have averaged the incident ionizing flux across the transverse directions. The first expression follows from $\lambda_{\nu}$ being constant across the cell by definition. [^4] The astute reader may wonder if this assumption contradicts our earlier observation that $n_{\rm HI}$ is not constant within cells. It does not because the isotropically-averaged opacity over some in-homogeneous region of the IGM still captures the average effect of the density fluctuations. The assumption here is that all information about the un-resolved in-homogeneity in $n_{\rm HI}$ is encoded in $\lambda_{\nu}$ (via whatever function is used for Eq. 4.11). Next, we average $\Gamma_{\rm HI}^{ij}$ in the incident direction across the length of the cell, $\Delta x_{\rm cell}$, and sum over all rays. The path length traversed by the ray is $\Delta s_{ij}$, which is always $\leq\Delta x_{\rm cell}$ (as enforced by our choice of time step, $\Delta t=\Delta x_{\rm cell}/c$). This gives
@@
In the second equality, we have evaluated the integral and used the fact that $A_{\rm cell}\Delta x_{\rm cell}=V_{\rm cell}$. Importantly, we see that Eq. 4.13 simplifies to Eq. 4.7 in the limit that $\lambda_{\nu}=1/[\sigma_{\rm HI}^{\nu}n_{\rm HI}]$, which is the case in Standard mode. This demonstrates that Eq. 4.13 does not depend on the plane-parallel geometry assumed in its derivation. Eq. 4.11 and 4.13 are updated iteratively until a self-consistent solution for $\lambda$ and $\Gamma_{\rm HI}$ is reached.
@@
Eq. 4.10-4.13 apply to highly ionized regions. To track the growth of these regions, we assume that the the boundary between highly ionized and fully neutral gas (that is, the I-front) is infinitely thin - this is the so-called “moving screen” approximation [^5] Note that a similar assumption is used in the RadHydro code. . This is equivalent to assuming that the I-front width, $\Delta x_{\rm IF}$, is always much less than $\Delta x_{\rm cell}$. Cosmological I-fronts are generally $10$s of pkpc across, with widths that vary depending on density and the spectrum of incident radiation [82]. Given our earlier discussion of the intended application of FlexRT’s Generalized Opacity formalism, it is reasonable to expect that this condition should hold in most relevant contexts (that is, when RT cells are $\sim$ a few Mpc). In subsequent sections, we will highlight this assumption whenever it becomes relevant (as it will in several instances). We will also explictly demonstrate the regime in which this approximation is valid in §6. The evolution of the ionized fraction $x_{ion}$ in a cell is (in the non-relativistic limit)
$\frac{dx_{\rm ion}}{dt}=\frac{v_{\rm IF}}{\Delta x_{\rm cell}}=\frac{F_{\gamma}}{(1+\chi)n_{\rm H}\Delta x_{\rm cell}}$ (4.14)

### v2

@@ 1 Introduction
[…]
@@
On the other hand, the fully numerical approach is well-suited for studying in detail the state and physics of the reionizing IGM. In recent years, this approach has generally moved in the direction of fully coupled radiative hydrodynamics simulations, many of which also simulate the galaxies, including star formation, AGN, and feedback processes [e.g. 49, 51, 55, 56, 52]. With this comes the great advantage that the simulations can now leverage the constraints provided by direct observations of the galaxy population. But the state-of-the-art simulations can cost millions of core hours per run, allowing for only a handful of simulations to be run per year, which precludes parameter space studies. The limited number of such simulations available also makes it difficult to disentangle the effects of approximations and differing sub-grid modeling choices, which are required in even the most state-of-the-art simulations [e.g. 57, 58]. It is also worth noting that the moment-based RT methods that are most widely used today truncate the moment hierarchy at the dipole moment of the intensity field by making an ansatz for the quadrupole, and therefore start from an approximate form of the RT equation. Hence, these methods also do not converge to the exact RT solution in any limit of the parameters related to numerical implementation (e.g. number of resolution elements).
@@ 2.1.1 Moment-Based RT methods
The high dimensionality of the RT equation makes it one of the most computationally intensive equations to solve numerically, even in relatively simple cases. A common approach to reducing the dimensionality is to take moments of the RT equation in the angular dimensions, and then truncate the hierarchy at the dipole level with an ansatz for the quadrupole moment, i.e. the Eddington tensor. The two most prominent ansatzes in use for reionization codes today are the M1 closure [61] and the optically thin Eddington tensor [OTVET; 62]. The equations governing the first two moments – the radiation energy density and flux – take on the form of conservation laws, analogous to the Euler equations of fluid dynamics, for which very efficient numerical solvers have been developed. Such codes are also relatively straightforward to parallelize across many CPUs because of their fluid-like description of the radiation field. Indeed, the M1 closure has the particularly convenient property that the Eddington tensor becomes a local function of radiation energy density and flux. Owing to its computational efficiency, the moment approach has become the most widely adopted type of RT algorithm used in almost all of the recent radiation hydrodynamics simulations of reionization, many of which include galaxy formation physics [49, 50, 55, 51, 52]. Indeed, without the moment-based RT implementations, the unprecedented scales and resolutions achieved by these simulations would not be possible.
The accuracy of moment-based codes has been quantified for a suite of idealized test problems in the Cosmological Radiative Transfer Comparison Project [59, 63], and in other works [e.g. 48, 64]. To our knowledge, however, there has been no quantitative study on the inaccuracies introduced by the fluid-like approximation to the radiation field, which arises from the local approximation to the Eddington tensor in the M1 method, in the cosmological reionization problem. There is good motivation for such a study. Ref. [65] considered several idealized RT problems that are relevant for reionization and found that moment-based methods tend to: (1) over-ionize dense, self-shielding absorption systems, which can result in a factor of $\sim 2$ under-prediction of the photoionization rate of intergalactic gas for a given emissivity; and (2) err in the structure of the ionizing radiation background on scales smaller than the mean free path (MFP). For instance, they found that the M1 closure significantly underestimates the fluctuations on these scales. This effect is visualized nicely in Figure 5 of Ref. [23], who noted that RT simulations run with Aton broadly exhibited a more uniform ionizing background compared to models generated with EX-CITE, which is essentially a ray tracing algorithm.
@@ 3.1 Ray creation and propagation
In traditional time-independent ray tracing schemes, rays would be traced away from the source until they either terminate in a neutral region or all their photons are absorbed by the IGM [76, 46, 23]. In FlexRT, “rays” are treated instead like photon packets that travel a distance $\Delta x_{\rm RT}$, the RT cell size, during each time step. Specifically, a ray starting a time step at position $(x_{0},y_{0},z_{0})$ travels to $(x_{1},y_{1},z_{1})=(x_{0},y_{0},z_{0})+\hat{v}\Delta x_{\rm RT}$ by the end of the time step, where $\hat{v}$ is the unit vector pointing along the original direction of the ray. Photons are absorbed into each cell intersected over this distance following the RT equation (see next section). Each ray retains its original unit vector assigned at casting until it is either split into child rays or merged with other rays (see next sub-sections). This approach, inspired by Monte-Carlo codes, naturally lends itself to the finite speed of light and “localizes” the rays, making FlexRT easier to parallelize (see §2.1.3). As we will see, this approach makes our procedure for merging of rays particularly straightforward.
To avoid numerical artifacts introduced by the HealPix structure, during each time step we randomly rotate the HealPix spheres used to cast rays around each source by a randomly chosen set of direction angles, $\alpha$, $\beta$, and $\gamma$. Each ray retains knowledge of these angles, so that the original HealPix unit vectors associated with it can be recovered when needed. We find that this procedure eliminates discernible structural artifacts even for simulations with just one or very few ionizing sources. Indeed, such artifacts can cause steady-state a-spherical structures in the radiation field around isotropic sources, and so removing them is necessary. An unavoidable side effect of this procedure is that the flux incident on a given cell will “flicker” from one time step to the next as the HealPix directions rotate, averaging to the correct value only over several time steps. This introduces artificial dependence on the ratio of the RT time step and the recombination timescale [^2] Note that all Monte Carlo RT (e.g. CRASH, Ref. [72]) codes are subject to this effect at some level, since they randomize the directions of photon packets emitted from sources. . Fortunately, as we will see, in most relevant applications, either the RT time step is much smaller than the recombination time scale or the local radiation field receives contributions from many sources in a way that averages out this flickering effect. We will comment on this point when it becomes relevant in subsequent sections.
@@ 3.3 Ray merging
Figure 1 shows a 2D representation of the steps in our merging scheme. Panel A shows an RT cell containing a set of rays. The lengths of the rays represent how many photons they contain, and rays with the same colors/line styles point in similar directions. The locations of the bases of the arrows denote the $(x,y)$ positions of the rays. In panel B, we throw down a randomly rotated “pizza” (analogous to a HealPix sphere in 3D) and assign each ray to the slice it occupies. Note that we have consolidated all rays to the center of the panel (ignoring for now their spatial positions) to make it easy to compare their directions. In panel C, rays within the same slice are merged. Note that two of the slices are empty and one of them has only a single ray, so no merging occurs in those slices. In the slice with the red rays, the longest ray had enough photons to be exempt from merging, and the other two were combined. Panel D shows again $(x,y)$ positions of the merged rays, which are the photon-weighted averaged positions of the constituent rays (see again panel A). Note that rays that were not merged with any others (long solid red and dotted black) have not changed their positions.
Figure 1: 2D representation of how ray merging works in FlexRT. Panel A shows a cell containing several rays. The lengths of the arrows represent the number of photons in each ray, and rays with same color/line style point in similar directions. The location of the bases of the arrows denotes the $(x,y)$ positions of the rays. In panel B, we throw down a randomly rotated “pizza” (analogous to a HealPix sphere in 3D) and group rays by the “slice” that they occupy, placing their bases in the center of the panel to aid comparison of their directions. We merge rays occupying the same slice in panel C, excluding the longest (red) ray, since it has enough photons to be exempt. Merged rays point along the bisector of the slice they occupy (i.e. the HealPix pixel direction). In panel D, we show that the positions of merged rays are the photon-weighted average of their constituents.
@@ 4.1 Direct solution of the chemistry equations (“Standard Mode”)
In this mode, we assume that all small-scale fluctuations in the HI number density ($n_{\rm HI}$) are fully resolved by our simulation. That is, $n_{\rm HI}$ does not fluctuate below the scale of the RT cell, $\Delta x_{\rm RT}$. Under this assumption, we can write an initial guess for the HI photo-ionization rate $\Gamma_{\rm HI}$ at the beginning of an RT time step of length $\Delta t=\Delta x_{\rm RT}/c$ as:
@@
where $T$ is temperature, $C_{\rm HI}(T)$ is the collisional ionization coefficient of HI, $\alpha(T)$ is the recombination coefficient of HII, $n_{e}$ is the electron number density and $n_{\rm HII}$ is the HII number density. Updating Eq. 4.2 using Eq. 4.1 would ignore the fact that $n_{\rm HI}$ and $\Gamma_{\rm HI}$ may evolve during $\Delta t$. Following Ref. [46], we account for this by solving for the time-dependence of the ionized fraction $x_{i}$ during $\Delta t$ assuming that the recombination rate remains constant (Eq. 12-14 of Ref. [46]),
@@
We iterate Eq. 4.3-4.7 several times to achieve convergence. Ref. [46] pointed out that this procedure gives the exact solution for $n_{\rm HI}$, no matter how large $\Delta t$ is, if recombinations are neglected. Generally, $n_{\rm HI}$ evolves most quickly inside ionization fronts, which are generally very thin ($10$s of ckpc [77]) and move very fast ($10^{3}-10^{4}$ km/s), such that the integrated number of recombinations taking place inside I-fronts is small compared to the number in ionized gas elsewhere. Thus, in the situations where the iteration procedure described above is most important, it is generally a good approximation to neglect recombinations. Note that the “inside” of the I-front refers to gas within the thin boundary between highly ionized ($x_{\rm HI}\sim 0$) and fully neutral ($x_{\rm HI}\sim 1$) gas.
However, one problem with Eq. 4.7 is that if $\Delta x_{\rm RT}$ is much larger than the I-front width $\Delta x_{\rm IF}$, the I-front will be “smeared out” uniformly over the cell containing it. This is because $n_{\rm HI}$ is assumed to have a uniform value everywhere in the cell. A more accurate treatment would divide the cell into a highly ionized and fully neutral component, and track the fraction of the cell in each state as the I-front crosses it. As we will see in the next sub-section, this is precisely the approach used in the Generalized Opacity Mode of FlexRT. In Standard Mode, the error resulting from this smearing effect is that the recombination rate in partially ionized cells is too low by approximately a factor of $\langle x_{i}\rangle$ whenever $\Delta x_{\rm RT}>>\Delta x_{\rm IF}$. As such, Standard Mode can under-estimate the integrated recombination rate if $\Delta x_{\rm RT}$ is large enough - we will return to this point in §6.
@@
In the first term on the RHS, $\mathcal{H}$ and $\Lambda$, are the net radiative heating and cooling rates, respectively, $n_{\rm tot}$ is the total number density of gas particles (neutrals, ions, and electrons), and $k_{\rm B}$ is Boltzmann’s constant. We include heating by photo-ionizations as well as cooling by recombinations, collisional ionizations and excitations, free-free interactions, and Compton scattering off the CMB. The second term, wherein $H(z)$ is the Hubble parameter, accounts for adiabatic cooling by cosmic expansion. The third term accounts for changes in temperature from an increase or decrease in the number of particles in the gas. The last term accounts for heating/cooling by compression/expansion, and is proportional to the change in the density in units of the mean, $\Delta$.
@@ 4.2 General treatment (“Generalized Opacity Mode”)
FlexRT’s second mode is built around relaxing the assumption that small-scale fluctuations in $n_{\rm HI}$ are fully resolved. Relaxing this assumption allows for each cell to have some un-resolved spatial distribution of $n_{\rm HI}$ inside it that would not be accounted for in Standard Mode. In this case, the equations in the previous section become ill-defined, since they rely on assuming a single value for $n_{\rm HI}$ in each cell. A commonly-employed method to correct for the effect of un-resolved density fluctuations is to boost the recombination rate (and thus, the residual $n_{\rm HI}$ in photo-ionization equilibrium) in each cell by a sub-grid clumping factor [e.g. 38]. However, this is not the most general approach, since it assumes that the ionizing opacity is determined uniquely by the recombination rate (e.g. Eq. 17 of Ref. [79]). This is not true in the presence of dense, self-shielding structures that harbor neutral gas, wherein the assumption of photo-ionization equilibrium may not hold. Specifically, neutral gas embedded in dense clumps may be in the process of photo-evaporation, or low-density ionized gas may be accreting onto denser structures and rapidly recombining. In these scenarios, the ionization rate can be larger or smaller than the recombination rate, respectively.
@@
$\tau_{\nu}\equiv\kappa_{\nu}\Delta s\equiv\Delta x_{\rm RT}/\lambda_{\nu}$ (4.9)
@@
Our final task is to provide a general expression for $\Gamma_{\rm HI}$ given knowledge of $\kappa_{\nu}$. Indeed, if $\Delta x_{\rm RT}$ were assumed to always be $<<\lambda_{\nu}$, Eq. 4.10 would be sufficient. However, as already mentioned, a key feature of FlexRT is that it should admit the use of large (up to several Mpc) RT cells without substantial loss of accuracy. As such, we must allow for scenarios in which $\lambda_{\nu}\lesssim\Delta x_{\rm RT}$. We can do this as follows. First, we assume (without loss of generality, as we will see) that all rays $j$ during the time step $\Delta t$ are incident on one face of cell $i$ traveling in the direction normal to that face. We can write the contribution of ray $j$ to $\Gamma_{\rm HI}$ at a distance $x_{j}$ into the cell as
@@
where $A_{\rm cell}\equiv\Delta x_{\rm RT}^{2}$ and we have averaged the incident ionizing flux across the transverse directions. The first expression follows from $\lambda_{\nu}$ being constant across the cell by definition. [^4] The astute reader may wonder if this assumption contradicts our earlier observation that $n_{\rm HI}$ is not constant within cells. It does not because the isotropically-averaged opacity over some in-homogeneous region of the IGM still captures the average effect of the density fluctuations. The assumption here is that all information about the un-resolved in-homogeneity in $n_{\rm HI}$ is encoded in $\lambda_{\nu}$ (via whatever function is used for Eq. 4.11). Next, we average $\Gamma_{\rm HI}^{ij}$ in the incident direction across the length of the cell, $\Delta x_{\rm RT}$, and sum over all rays. The path length traversed by the ray is $\Delta s_{ij}$, which is always $\leq\Delta x_{\rm RT}$ (as enforced by our choice of time step, $\Delta t=\Delta x_{\rm RT}/c$). This gives
@@
In the second equality, we have evaluated the integral and used the fact that $A_{\rm cell}\Delta x_{\rm RT}=V_{\rm cell}$. Importantly, we see that Eq. 4.13 simplifies to Eq. 4.7 in the limit that $\lambda_{\nu}=1/[\sigma_{\rm HI}^{\nu}n_{\rm HI}]$, which is the case in Standard mode. This demonstrates that Eq. 4.13 does not depend on the plane-parallel geometry assumed in its derivation. Eq. 4.11 and 4.13 are updated iteratively until a self-consistent solution for $\lambda$ and $\Gamma_{\rm HI}$ is reached.
@@
Eq. 4.10-4.13 apply to highly ionized regions. To track the growth of these regions, we assume that the boundary between highly ionized and fully neutral gas (that is, the I-front) is infinitely thin - this is the so-called “moving screen” approximation [^5] Note that a similar assumption is used in the RadHydro code. . This is equivalent to assuming that $\Delta x_{\rm IF}<<\Delta x_{\rm RT}$. Indeed, it is the opposite assumption to that made in the §4.1, namely that $n_{\rm HI}$ is uniform within partially ionized cells. Given our earlier discussion of the intended application of FlexRT’s Generalized Opacity formalism, it is reasonable to expect that this condition should hold in most relevant contexts (that is, when RT cells are $\sim$ a few Mpc). In subsequent sections, we will highlight this assumption whenever it becomes relevant (as it will in several instances). We will also explictly demonstrate the regime in which this approximation is valid in §6. The evolution of the ionized fraction $x_{ion}$ in a cell is (in the non-relativistic limit)
$\frac{dx_{\rm ion}}{dt}=\frac{v_{\rm IF}}{\Delta x_{\rm RT}}=\frac{F_{\gamma}}{(1+\chi)n_{\rm H}\Delta x_{\rm RT}}$ (4.14)